# Stable dark matter from Pauli blocking in the degenerate fermion background with Quantum Field Theory

### v1

@@ 4 A Specific Model
$\begin{split}\phi(t)=\bar{\phi}\left(\frac{a(t)}{a(t_{0})}\right)^{-3/2}\cos(m_{\phi}t).\end{split}$ (4.1)
@@
$\begin{split}\rho_{\phi}(t)=\rho_{\phi}(t_{0})\left(\frac{a(t)}{a(t_{0})}\right)^{-3}\simeq\rho_{\phi}(t_{0})\left(\frac{t}{t_{0}}\right)^{-2},\end{split}$ (4.2)
@@
$\begin{split}\rho_{\phi}(t_{0})=\frac{1}{2}m_{\phi}^{2}\bar{\phi}^{2}\simeq 3.7\times 10^{-47}\,{\rm GeV}^{4},\end{split}$ (4.3)
@@
$\begin{split}n_{\phi}(t)=\frac{\rho_{\phi}(t)}{m_{\phi}}=3.7\times 10^{-35}\,{\rm GeV}^{3}\left(\frac{10^{-3}\,{\rm eV}}{m_{\phi}}\right)\left(\frac{t}{t_{0}}\right)^{-2}.\end{split}$ (4.4)
For a fixed DM energy density in Eq. (4.3), the variable $q$ can be written as
$\begin{split}q\equiv\frac{\lambda\bar{\phi}}{m_{\phi}}=8.6\times 10^{-12}\left(\frac{\lambda}{10^{-12}}\right)\left(\frac{10^{-3}\,{\rm eV}}{m_{\phi}}\right)^{2}.\end{split}$ (4.5)
The lifetime without Pauli blocking in Eq. (2.4) is
$\begin{split}t_{\phi}\simeq 1.6\times 10^{13}\sec\left(\frac{10^{-12}}{\lambda}\right)^{2}\left(\frac{10^{-3}\,{\rm eV}}{m_{\phi}}\right),\end{split}$ (4.6)
however this is extended due to the Pauli blocking as in Eq. (2.34) to
$\begin{split}t_{\phi,ext}\simeq 132\times t_{0}\left(\frac{10^{-3}\,{\rm eV}}{m_{\phi}}\right)^{2},\end{split}$ (4.7)
which is independent of the coupling $\lambda$, once $t_{\phi,ext}>t_{\phi}$.
@@
$\begin{split}n_{\psi}=\frac{p_{\psi}^{3}}{6\pi^{2}},\qquad\rho_{\psi}=\frac{m_{\phi}^{4}}{128\pi^{2}}.\end{split}$ (4.8)
@@
$\begin{split}\frac{d}{dt}(\rho_{\phi}+\rho_{\psi})+3H\rho_{\phi}+4H\rho_{\psi}=0,\end{split}$ (4.9)
@@
$\begin{split}\rho_{\phi}(t)=-\frac{m_{\phi}^{4}}{96\pi^{2}}+\left(\rho_{\phi}(t_{i})+\frac{m_{\phi}^{4}}{96\pi^{2}}\right)\left(\frac{a(t)}{a(t_{i})}\right)^{-3},\end{split}$ (4.10)
@@
$\mathcal{L}_{\text{Yukawa }}\supset-\frac{1}{\sqrt{2}}Y_{\Delta}^{ij}\overline{l_{L}^{iC}}\cdot\Delta l_{L}^{j}+h.c.$ (4.11)
After the triplet Higgs field develops a VEV $v_{\Delta}$, left-handed neutrino masses are generated as
$\left(\mathcal{M}_{\nu}\right)_{ij}=Y_{\Delta}^{ij}v_{\Delta}.$ (4.12)
@@
$V=$ $V_{1}+V_{2}$ (4.13)
$V_{1}=$ $-\hat{\mu}_{1}{}^{2}\left|\Phi_{1}\right|^{2}+\frac{1}{2}\lambda_{1}\left|\Phi_{1}\right|^{4}$
$+\hat{\mu}_{3}{}^{2}\operatorname{Tr}\left(\Delta^{\dagger}\Delta\right)+\frac{1}{2}\Lambda_{1}\left(\operatorname{Tr}\left(\Delta^{\dagger}\Delta\right)\right)^{2}+\frac{1}{2}\Lambda_{2}\left(\left(\operatorname{Tr}\left(\Delta^{\dagger}\Delta\right)\right)^{2}-\operatorname{Tr}\left[\left(\Delta^{\dagger}\Delta\right)^{2}\right]\right)$
@@
where $\Phi_{1}$ is the SM doublet Higgs and $S$ is the real singlet scalar of dark matter candidate.
The neutral component of $S,\Phi_{1}$ and $\Delta$ are denoted by $s,\phi^{0}$ and $\delta^{0}$, respectively. The mass squared matrix
$M_{ij}^{2}\quad(i,j)\quad\mathrm{run}\quad\left(s,\phi^{0},\delta^{0}\right)$ (4.16)
@@
$U=\left(\begin{array}[]{ccc}U_{s1}&U_{s2}&U_{s3}\\ U_{\phi^{0}1}&U_{\phi^{0}2}&U_{\phi^{0}3}\\ U_{\delta^{0}1}&U_{\delta^{0}2}&U_{\delta^{0}3}\end{array}\right).$ (4.17)
The DM $\phi_{1}$ decays into two neutrino pair through the mixing $U_{\delta^{0}1}$ and the Yukawa coupling (4.11). The decay rate is
@@
$\simeq 8\times 10^{-43}\,{\rm GeV}\left(\frac{m_{\phi_{1}}}{10^{-3}\,{\rm eV}}\right)\left(\frac{m_{\nu}}{10^{-6}\,{\rm eV}}\right)^{2}\left(\frac{1\,{\rm keV}}{v_{\Delta}}\right)^{2}\left(\frac{\absolutevalue{U_{\delta^{0}1}}^{2}}{10^{-11}}\right),$ (4.18)
@@
$U_{\delta^{0}1}\sim$ $\frac{M_{\delta^{0}}^{2}}{M_{\delta^{0}\delta^{0}}^{2}}\simeq\frac{\lambda_{6}v^{2}}{\hat{\mu}_{3}^{2}}\simeq 10^{-5}\left(\frac{\lambda_{6}}{0.1}\right)\left(\frac{10^{4}\mathrm{GeV}}{\hat{\mu}_{3}}\right)^{2}.$ (4.19)
Though we normalized the decay rate $\Gamma$ in Eq. (4.18) by about the age of Universe $t_{0}$ and the actual bound is $\Gamma^{-1}>\mathcal{O}(10)t_{0}$ [1, 2, 3], as we have seen above that smaller $t_{\phi}$ is enough to have long enough $t_{\phi,ext}$. In fact, a much larger element (4.19) is allowed. Since the neutrinos are in the weak $SU(2)$ doublet, in fact, DM $\phi$ can decay into two photons through loop processes involving neutrinos, the $W$ boson and electrons. This decay rate is very suppressed to be small enough [20].
@@ 5 Conclusion
Acknowledgments. W.Cho, K.-Y. Choi, and J.Joh acknowledge the financial support from National Research Foundation(NRF) grant funded by the Korea government (MEST) NRF-2022R1A2C1005050. This work of OS was supported in part by JSPS KAKENHI Grant Number 23K03402.

### v2

@@ 4 A Specific Model
$\begin{split}\phi(t)=\bar{\phi}\left(\frac{a(t)}{a(t_{0})}\right)^{-3/2}\cos(m_{\phi}t).\end{split}$ (79)
@@
$\begin{split}\rho_{\phi}(t)=\rho_{\phi}(t_{0})\left(\frac{a(t)}{a(t_{0})}\right)^{-3}\simeq\rho_{\phi}(t_{0})\left(\frac{t}{t_{0}}\right)^{-2},\end{split}$ (80)
@@
$\begin{split}\rho_{\phi}(t_{0})=\frac{1}{2}m_{\phi}^{2}\bar{\phi}^{2}\simeq 3.7\times 10^{-47}\,{\rm GeV}^{4},\end{split}$ (81)
@@
$\begin{split}n_{\phi}(t)=\frac{\rho_{\phi}(t)}{m_{\phi}}=3.7\times 10^{-35}\,{\rm GeV}^{3}\left(\frac{10^{-3}\,{\rm eV}}{m_{\phi}}\right)\left(\frac{t}{t_{0}}\right)^{-2}.\end{split}$ (82)
For a fixed DM energy density in Eq. (81), the variable $q$ can be written as
$\begin{split}q\equiv\frac{\lambda\bar{\phi}}{m_{\phi}}=8.6\times 10^{-12}\left(\frac{\lambda}{10^{-12}}\right)\left(\frac{10^{-3}\,{\rm eV}}{m_{\phi}}\right)^{2}.\end{split}$ (83)
The lifetime without Pauli blocking in Eq. (4) is
$\begin{split}\tau_{\phi,0}\simeq 1.6\times 10^{13}\sec\left(\frac{10^{-12}}{\lambda}\right)^{2}\left(\frac{10^{-3}\,{\rm eV}}{m_{\phi}}\right),\end{split}$ (84)
however this lifetime is extended due to the Pauli blocking.
As explained in Eq. (33), the scalar $\phi$ would decay quickly when the number density becomes smaller than that in the Fermi sphere, $n_{\phi}(t)\lesssim\frac{p_{\psi}^{3}}{6\pi^{2}}$, when the Pauli blocking is inefficient. Using this equation and Eq. (82) with $p_{\psi}\simeq m_{\phi}/2$, we find the extended lifetime as
$\begin{split}\tau_{\phi}\simeq 132\times t_{0}\left(\frac{10^{-3}\,{\rm eV}}{m_{\phi}}\right)^{2},\end{split}$ (85)
which is independent of the coupling $\lambda$, once $\tau_{\phi}>\tau_{\phi,0}$.
There is also a constraint on the mass of $\phi$, $m_{\phi}$, for Pauli blocking to work. To fill the Fermi sphere with Fermi momentum $p_{\psi}$, we need at least a number density larger than the number density in the Fermi sphere as
$\begin{split}n_{\phi}\gg\frac{p_{\psi}^{3}}{6\pi^{2}}.\end{split}$ (86)
Using Eq. (81) and Eq. (82) with $p_{\psi}\simeq m_{\phi}/2$ for $m_{\psi}\ll m_{\phi}$, we find the upper bound on the mass
$\begin{split}m_{\phi}\lesssim(16\pi^{2}\rho_{\rm DM})^{1/4}\simeq 0.01\,{\rm eV}.\end{split}$ (87)
Therefore for larger value of $m_{\phi}$ than this, the Pauli blocking does not work due to the insufficient number density of $\phi$. Note that, for subdominant DM component, the constraint becomes stronger since its energy density is smaller by the fractional factor.
In this paper, we consider that $\phi$ and $\psi$ do not have further scattering between each other except the decay and inverse decay, since the scatterings may redistribute the momentum spectrum of $\psi$ and destroy the Pauli blocking. For this, we need small Yukawa coupling, so that the scattering rate $n\langle\sigma v\rangle\sim\lambda^{4}\rho_{\phi}/m_{\phi}^{3}$ is much smaller than the Hubble rate $H$ at present. This gives the upper bound on the Yukawa coupling as
$\begin{split}\lambda\ll\left(\frac{Hm_{\phi}^{3}}{\rho_{\phi}}\right)^{1/4}\simeq 2\times 10^{-8}\left(\frac{m_{\phi}}{\rm 10^{-3}eV}\right)^{3/4}.\end{split}$ (88)
For this parameter, the Pauli blocking is not affected in our model.
For very small Yukawa coupling and mass, the lifetime in the vacuum becomes longer than the age of the Universe, then the Pauli blocking becomes meaningless at the present Universe.
@@
$\begin{split}n_{\psi}=\frac{p_{\psi}^{3}}{6\pi^{2}},\qquad\rho_{\psi}=\frac{m_{\phi}^{4}}{128\pi^{2}}.\end{split}$ (89)
@@
$\begin{split}\frac{d}{dt}(\rho_{\phi}+\rho_{\psi})+3H\rho_{\phi}+4H\rho_{\psi}=0,\end{split}$ (90)
@@
$\begin{split}\rho_{\phi}(t)=-\frac{m_{\phi}^{4}}{96\pi^{2}}+\left(\rho_{\phi}(t_{i})+\frac{m_{\phi}^{4}}{96\pi^{2}}\right)\left(\frac{a(t)}{a(t_{i})}\right)^{-3},\end{split}$ (91)
@@
$\mathcal{L}_{\text{Yukawa }}\supset-\frac{1}{\sqrt{2}}Y_{\Delta}^{ij}\overline{l_{L}^{iC}}\cdot\Delta l_{L}^{j}+h.c.$ (92)
After the triplet Higgs field develops the vacuum expectation value (VEV) $v_{\Delta}$, left-handed neutrino masses are generated as
$\left(\mathcal{M}_{\nu}\right)_{ij}=Y_{\Delta}^{ij}v_{\Delta}.$ (93)
@@
$V=$ $V_{1}+V_{2}$ (94)
$V_{1}=$ $-\mu_{1}{}^{2}\left|\Phi\right|^{2}+\frac{1}{2}\lambda_{1}\left|\Phi\right|^{4}$
$+\mu_{3}{}^{2}\operatorname{Tr}\left(\Delta^{\dagger}\Delta\right)+\frac{1}{2}\Lambda_{1}\left(\operatorname{Tr}\left(\Delta^{\dagger}\Delta\right)\right)^{2}+\frac{1}{2}\Lambda_{2}\left(\left(\operatorname{Tr}\left(\Delta^{\dagger}\Delta\right)\right)^{2}-\operatorname{Tr}\left[\left(\Delta^{\dagger}\Delta\right)^{2}\right]\right)$
@@
where $\Phi$ is the SM doublet Higgs and $S$ is the real singlet scalar of dark matter candidate. Around the VEVs of $\Delta$, $\Phi$ and $S$, which are $v_{\Delta},v\simeq 246$ GeV and $v_{s}$ respectively, the real neutral components of those fields are expanded as
$\Delta=\left(\begin{array}[]{cc}0&,0\\ \frac{1}{\sqrt{2}}(v_{\Delta}+\delta^{0})&,0\end{array}\right),$ (99)
$\Phi=\left(\begin{array}[]{c}0\\ \frac{1}{\sqrt{2}}(v+\phi^{0})\end{array}\right),$ (102)
$S=v_{s}+s,$ (103)
and VEVs are expressed as
$v_{\Delta}$ $\simeq-\frac{\lambda_{S1}\Lambda_{6}v^{2}}{v^{2}\left(\lambda_{S1}(\Lambda_{4}-\Lambda_{5})-2\lambda_{6}^{2}\right)+2\lambda_{S1}\mu_{3}^{2}},$ (104)
$v_{s}$ $\simeq-\frac{\lambda_{6}}{\lambda_{S1}}v_{\Delta}.$ (105)
The mass squared matrix
$M_{ij}^{2},\quad\mathrm{where}\quad(i,j)\quad\mathrm{run}\quad\left(s,\phi^{0},\delta^{0}\right)$ (106)
@@
$U=\left(\begin{array}[]{ccc}U_{s1}&U_{s2}&U_{s3}\\ U_{\phi^{0}1}&U_{\phi^{0}2}&U_{\phi^{0}3}\\ U_{\delta^{0}1}&U_{\delta^{0}2}&U_{\delta^{0}3}\end{array}\right).$ (107)
The DM $\phi_{1}$ decays into two neutrino pair through the mixing $U_{\delta^{0}1}$ and the Yukawa coupling (92). The decay rate is
@@
$\simeq 8\times 10^{-43}\,{\rm GeV}\left(\frac{m_{\phi_{1}}}{10^{-3}\,{\rm eV}}\right)\left(\frac{m_{\nu}}{10^{-6}\,{\rm eV}}\right)^{2}\left(\frac{1\,{\rm keV}}{v_{\Delta}}\right)^{2}\left(\frac{\absolutevalue{U_{\delta^{0}1}}^{2}}{10^{-11}}\right),$ (108)
@@
$U_{\delta^{0}1}\sim$ $\frac{M_{s\delta^{0}}^{2}}{M_{\delta^{0}\delta^{0}}^{2}}\simeq\frac{\lambda_{6}v^{2}}{{\mu}_{3}^{2}}\simeq 10^{-5}\left(\frac{\lambda_{6}}{0.1}\right)\left(\frac{10^{4}\mathrm{GeV}}{\mu_{3}}\right)^{2}.$ (109)
where we have assumed $\mu_{3}^{2}\gg v^{2}$. Though we normalized the decay rate $\Gamma$ in Eq. (108) by about the age of Universe $t_{0}$ and the actual bound is $\Gamma^{-1}>\mathcal{O}(10)t_{0}$ [1, 2, 3], as we have seen above that smaller $t_{\phi}$ is enough to have long enough $t_{\phi,ext}$. In fact, a much larger element (109) is allowed. Since the neutrinos are in the weak $SU(2)$ doublet, in fact, DM $\phi$ can decay into two photons through loop processes involving neutrinos, the $W$ boson and electrons. This decay rate is very suppressed to be small enough [22].
The neutrinos produced from the decay of DM occupies the Fermi surface inside the phase space with the momentum of $p_{\psi}=m_{\phi_{1}}/2$. The number density of this neutrino corresponds to
$\begin{split}n_{\nu}=\frac{p_{\psi}^{3}}{6\pi^{2}}=2.7\times 10^{5}{\rm cm}^{-3}\left(\frac{m_{\phi}}{0.01\,{\rm eV}}\right)^{3}.\end{split}$ (110)
This is comparable to the cosmic background neutrino in the standard cosmological model which is
$\begin{split}n_{\nu,\rm SM}=56{\rm cm}^{-3},\end{split}$ (111)
per flavor. The enhanced number density might be probed in the future experiments of the cosmic neutrino background like PTOLEMY-like experiment [23].
@@ 5 Conclusion
Acknowledgements
W.Cho, K.-Y. Choi, and J.Joh acknowledge the financial support from National Research Foundation(NRF) grant funded by the Korea government (MEST) NRF-2022R1A2C1005050. This work of OS was supported in part by JSPS KAKENHI Grant Number 23K03402.